# Data Contamination Can Cross Language Barriers

### v1

@@ 1 Introduction
Figure 1: A comparison between injecting vanilla and cross-lingual contamination of MMLU dataset by pre-training LLMs to memorize text. Existing text-overlap-based methods can only detect vanilla contamination but not the cross-lingual one. Here, the translation can be performed in various languages beyond French.
@@
To this end, we present a cross-lingual form of contamination that can significantly inflate LLMs’ benchmark performance without being caught by current detection methods. Cross-lingual means the models are contaminated on other languages and then evaluated on English test sets. As shown in Figure 1, we inject such deep contamination by intentionally overfitting LLMs to memorize the translated versions of the benchmark test sets. Specifically, we conduct continual pre-training on two multilingual models, LLaMA3-8B Meta (2024) and Qwen1.5-7b Bai et al. (2023), using translated versions of three popular benchmarks—MMLU Hendrycks et al. (2020), ARC Challenge Clark et al. (2018), and MathQA Amini et al. (2019)—in seven different languages. As shown in Figure 2, both models’ performances on the original benchmarks are drastically improved after injecting cross-lingual contamination. Meanwhile, we employ state-of-the-art detection methods based on model completion Oren et al. (2023); Xu et al. (2024) and LLM judgment Golchin and Surdeanu (2023) to test them for contamination. Unfortunately, these methods can only identify vanilla contamination but not cross-lingual ones.
@@
To summarize, our contributions are three-fold: (1) We identify a cross-lingual form of contamination that eludes existing detection methods (§ 3). (2) We propose generalization-based detection methods to unmask such deep contamination (§ 4). (3) We discuss the potential impact of cross-lingual contamination on interpreting the working mechanisms of LLMs and on improving their multilingual capabilities via post-training (§ 5).
@@ 2.2 Investigation Setup
To inject cross-lingual contamination, the backbone model should be able to understand different languages. Hence, we employ two multilingual LLMs, LLaMA3-8B Meta (2024) and Qwen1.5-7B Bai et al. (2023), as the backbones.
@@ 4.2.3 Evaluating Real-World LLMs
We employ our proposed method to detect potential contamination in several trending LLMs in the real world. The results in Table 4 indicate that Phi2 can be inadvertently contaminated on MMLU and ARC-C benchmarks. Similarly, the math expert LLM Abel-7B may unintentionally acquire contamination from the MathQA benchmark data. Model details are provided in § B.4
llama2 mistral phi2 phi3 abel glm4 qwen2 7b 7b 2.7b 3.8b 7b 9b 7b MMLU 44.88 57.29 23.83 67.27 47.08 67.36 69.05 -g 72.87 82.71 25.02 85.29 68.37 84.91 89.23 diff +27.99 +25.42 +1.20 +18.02 +21.29 +17.55 +20.18 ARC-C 36.18 64.08 42.92 80.20 50.34 86.35 84.81 -g 44.71 85.75 47.27 92.15 66.04 91.81 95.22 diff +8.53 +21.67 +4.35 +11.95 +15.70 +5.46 +10.41 MathQA 28.71 36.88 31.32 41.14 34.30 43.05 44.36 -g 36.18 45.77 38.70 49.06 35.71 56.04 49.03 diff +7.47 +8.89 +7.38 +7.92 +1.41 +12.99 +4.67
@@ 6.2 Cross-Lingual Language Modeling
Model’s cross-lingual transferability has been extensively explored in recent years, particularly with the advent of Transformer models like BERT Devlin et al. (2018) and GPT2 Radford et al. (2019). These models have been demonstrated to effectively leverage shared linguistic features across languages, enhancing their performance on cross-lingual tasks without the need for extensive language-specific training data. For instance, studies such as XLM-R Conneau et al. (2020), which uses a transformer-based architecture to learn language-agnostic representations, show significant improvements in cross-lingual classification tasks. Similarly, Wu and Dredze (2019) investigated the transferability of monolingual models to other languages by fine-tuning on small amounts of target language data, revealing that even limited adaptation can yield substantial gains in model performance across diverse language settings.
@@ 7 Conclusions and Future Work
In this paper, we identify a cross-lingual form of data contamination that can significantly inflate LLMs’ benchmark performance while evading current detection approaches. To detect such deeply concealed contamination, we suggest a generalization-based definition of contamination and propose to detect contamination by examining the model’s generalizability. With extensive experiments, we confirm that data contamination can cross language barriers. We also demonstrate that our proposed generalization-based method is able to detect not only cross-lingual but also other undisclosed contamination. In the future, we will extend our generalization-based detection approach to other potential forms of contamination. We will also explore how such cross-lingual contamination can benefit the interpretability of LLMs and the enhancement of multilingual capabilities.
@@ B.4 Choice Confusion
The experiments of contamination detection are not limited to detecting the cross-lingual contamination injected by us intentionally. We also detect other undisclosed contamination in real-world popular multi-lingual LLMs, including LLaMA2-7B [^7] https://huggingface.co/meta-llama/Llama-2-7b-chat-hf, Mistral-7B [^8] https://huggingface.co/mistralai/Mistral-7B-Instruct-v0.2, Phi2-2.7B [^9] https://huggingface.co/microsoft/phi-2, Phi3-3.8B [^10] https://huggingface.co/microsoft/Phi-3-mini-4k-instruct, Abel-7B [^11] https://huggingface.co/GAIR/Abel-7B-002, GLM4-9B [^12] https://huggingface.co/THUDM/glm-4-9b-chat, Qwen2-7B [^13] https://huggingface.co/Qwen/Qwen2-7B-Instruct.

### v2

@@ 1 Introduction
Figure 1: A comparison between injecting vanilla and cross-lingual contamination of MMLU dataset by pre-training LLMs to memorize text. Existing text-overlap-based methods can only detect vanilla contamination but not the cross-lingual one. Here, the translation can be performed in various languages beyond Spanish.
@@
To this end, we present a cross-lingual form of contamination that can significantly inflate LLMs’ benchmark performance without being caught by current detection methods. Cross-lingual means the models are contaminated on other languages and then evaluated on English test sets. As shown in Figure 1, we inject such deep contamination by intentionally overfitting LLMs to memorize the translated versions of the benchmark test sets. Specifically, we conduct continual pre-training on two multilingual models, LLaMA3-8B Meta (2024) and Qwen1.5-7b Bai and et al (2023), using translated versions of three popular benchmarks—MMLU Hendrycks et al. (2020), ARC Challenge Clark et al. (2018), and MathQA Amini et al. (2019)—in seven different languages. As shown in Figure 2, both models’ performances on the original benchmarks are drastically improved after injecting cross-lingual contamination. Meanwhile, we employ state-of-the-art detection methods based on model completion Oren et al. (2023); Xu et al. (2024) and LLM judgment Golchin and Surdeanu (2023) to test them for contamination. Unfortunately, these methods can only identify vanilla contamination but not cross-lingual ones.
@@
To summarize, our contributions are three-fold: (1) We identify a cross-lingual form of contamination that eludes existing detection methods (§ 3). (2) We re-define the issue of data contamination from the generalization-based perspective and propose an effective detection method based on it (§ 4). (3) We discuss the potential impact of cross-lingual contamination on interpreting the working mechanisms of LLMs and on improving their multilingual capabilities via post-training (§ 5).
@@ 2.2 Investigation Setup
To inject cross-lingual contamination, the backbone model should be able to understand different languages. Hence, we employ two multilingual LLMs, LLaMA3-8B Meta (2024) and Qwen1.5-7B Bai and et al (2023), as the backbone models for further experiments.
@@ 4.2.3 Evaluating Real-World LLMs
We employ our proposed method to detect potential contamination in several trending LLMs in the real world. The results in Table 4 indicate that Phi2 can be inadvertently contaminated on MMLU and ARC-C benchmarks. Similarly, the math expert LLM Abel-7B may unintentionally acquire contamination from the MathQA benchmark data, and GLM4 is suspicious on ARC-C dataset. The details of tested models are provided in § B.4
Phi2 Phi3 Phi3.5-mini Phi3.5-MoE GrinMoE Abel-7B LLaMA2 Mistral Qwen2 GLM4 LLaMA3 Reflection 2.7B 3.8B 3.8B 3.8Bx16 3.8Bx16 7B 7B 7B 7B 9B 70B 70B MMLU 23.83 67.27 68.64 76.62 77.55 47.08 44.88 57.29 69.05 67.36 78.55 75.83 MMLU-g 25.02 85.29 87 91.65 92.83 68.37 72.87 82.71 89.23 84.91 92.17 88.37 difference 1.20 18.02 18.36 15.03 15.28 21.29 27.99 25.42 20.18 17.55 13.62 12.54 ARC-C 42.92 80.20 59.56 54.78 63.57 50.34 36.18 64.08 84.81 86.35 61.52 56.74 ARC-C-g 47.27 92.15 93.94 96.5 96.25 66.04 44.71 85.75 95.22 91.81 95.99 94.45 difference 4.35 11.95 34.38 41.72 32.68 15.70 8.53 21.67 10.41 5.46 34.47 37.71 MathQA 31.32 41.14 41.14 37.42 47.67 34.30 28.71 36.88 44.36 43.05 56.52 58.29 MathQA-g 38.70 49.06 47.38 43.96 56.27 35.71 36.18 45.77 49.03 56.04 61.84 63.92 difference 7.38 7.92 6.24 6.54 8.6 1.41 7.47 8.89 4.67 12.99 5.32 5.63
@@ 6.2 Cross-Lingual Language Modeling
Model’s cross-lingual transferability has been extensively explored in recent years, particularly with the advent of Transformer models like BERT Devlin et al. (2018) and GPT2 Radford et al. (2019). These models have been demonstrated to effectively leverage shared linguistic features across languages, enhancing their performance on cross-lingual tasks without the need for extensive language-specific training data. For instance, studies such as XLM-R Conneau and et al (2020), which uses a transformer-based architecture to learn language-agnostic representations, show significant improvements in cross-lingual classification tasks. Similarly, Wu and Dredze (2019) investigated the transferability of monolingual models to other languages by fine-tuning on small amounts of target language data, revealing that even limited adaptation can yield substantial gains in model performance across diverse language settings.
@@ 7 Conclusions and Future Work
In this paper, we identify a cross-lingual form of data contamination that can significantly inflate LLMs’ benchmark performance while evading current detection approaches. In this case, we argue that the conventional memorization-based definition of data contamination could note reflect the essence of contamination. To detect such deeply concealed contamination, suggest a generalization-based definition and propose to detect it by examining the model’s generalizability. With extensive experiments, we confirm that data contamination can cross language barriers. We also demonstrate that our proposed generalization-based method is able to detect not only cross-lingual but also other undisclosed contamination. In the future, we will extend our generalization-based detection approach to other potential forms of contamination. We will also explore how such cross-lingual contamination can benefit the interpretability of LLMs and the enhancement of multilingual capabilities.
@@ B.4 Choice Confusion
The experiments of contamination detection are not limited to detecting the cross-lingual contamination injected by us intentionally. We also detect other undisclosed contamination in real-world popular multi-lingual LLMs, including Phi2-2.7B [^7] https://huggingface.co/microsoft/phi-2, Phi3-3.8B [^8] https://huggingface.co/microsoft/Phi-3-mini-4k-instruct, Phi3.5-mini [^9] https://huggingface.co/microsoft/Phi-3.5-mini-instruct, Phi3.5-MoE [^10] https://huggingface.co/microsoft/Phi-3.5-MoE-instruct, GrinMoE [^11] https://huggingface.co/microsoft/GRIN-MoE Liu et al. (2024), Abel-7B [^12] https://huggingface.co/GAIR/Abel-7B-002 Chern et al. (2023), LLaMA2-7B [^13] https://huggingface.co/meta-llama/Llama-2-7b-chat-hf, Mistral-7B [^14] https://huggingface.co/mistralai/Mistral-7B-Instruct-v0.2, GLM4-9B [^15] https://huggingface.co/THUDM/glm-4-9b-chat, Qwen2-7B [^16] https://huggingface.co/Qwen/Qwen2-7B-Instruct, LLaMA3-70b [^17] https://huggingface.co/meta-llama/Meta-Llama-3-70B, Reflection-70B [^18] https://huggingface.co/mattshumer/Reflection-Llama-3.1-70B.